# Beyond the non-Hermitian skin effect: scaling-controlled topology from Exceptional-Bound Bands

### v1

@@ I Introduction
In this work, however, we show that anomalous non-Hermitian scaling can occur without any skin effect at all. Departing from existing paradigms [1, 2, 4, 2, 19, 3, 20, 21, 7, 8], we propose a completely different mechanism for scale-induced topological transitions. Our approach relies on unique renormalization properties near an exceptional point (EP) due to eigenspace defectiveness. While EPs are already known for their extremely sensitive band dispersions [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32] and unconventional entanglement properties [33, 34, 35, 36, 37, 38, 39, 40], what has not been previously harnessed is the non-locality that emerges from projecting onto their defective eigenspaces.
To employ this non-locality for realizing novel scaling transitions, we specifically dimensionally extend these eigenspaces to form so-called exceptional-bound (EB) bands. These are entire energy bands protected by the presence of momentum-space singularities. As we shall show, EB bands possess unique algebraically decaying profiles that scale strongly with system size, harboring the propensity for scale-sensitive band engineering. Starting from a parent exceptional point, we demonstrate how we can build effective models with precisely scale-tunable couplings. This enables the same system to be topological at some system sizes but non-topological at other sizes, i.e., exhibiting scaling-controlled topological phase transitions. We develop a formalism for engineering such unprecedented scale-controlled transitions, fundamentally expanding the potential of non-Hermitian mechanisms in non-locally controlling entire band structures, beyond the reach of existing sensing application proposals.
@@ II.1 Unconventional scaling behavior from exceptional point defectiveness
$P(k)=\sum_{n\in occ.}|\psi_{n}(k)\rangle\langle\tilde{\psi}_{n}(k)|=\frac{1}{2}\left(\begin{array}[]{cc}1&-U(k)\\ -D(k)&1\end{array}\right),$ (2)
@@
Importantly, the off-diagonal correlator $U(k)=D(k)^{-1}\sim a_{0}k^{-B}$ diverges as $k\rightarrow 0$, since the projector should be ill-defined when the eigenvectors coalesce. Note that in order to satisfy $P^{2}=P$ in a non-Hermitian setting, this projector $P$ must be defined with respect to the biorthogonal basis which is given by $\mathcal{H}|\psi_{n}\rangle=\mathcal{E}|\psi_{n}\rangle$ and $\mathcal{H}^{\dagger}|\tilde{\psi}_{n}\rangle=\mathcal{E}^{*}|\tilde{\psi}_{n}\rangle$, with $\langle\tilde{\psi}_{n}|\psi_{n^{\prime}}\rangle=\delta_{nn^{\prime}}$.
@@
$=$ $\frac{1}{2}\left(\begin{array}[]{cc}R_{\Gamma}&-R_{\Gamma}UR_{\Gamma}\\ -R_{\Gamma}DR_{\Gamma}&R_{\Gamma}\end{array}\right).$ (12)
@@
$=$ $\frac{1}{2}\left(\begin{array}[]{cc}R_{\Gamma}UR_{\Gamma}DR_{\Gamma}-R_{\Gamma}&0\\ 0&R_{\Gamma}DR_{\Gamma}UR_{\Gamma}-R_{\Gamma}\end{array}\right)$ (18)
$=$ $-\frac{1}{2}R_{\Gamma}\left(\begin{array}[]{cc}UR^{c}_{\Gamma}U^{-1}&0\\ 0&U^{-1}R^{c}_{\Gamma}U\end{array}\right)R_{\Gamma}$ (21)
$=$ $-\frac{1}{2}\sum_{y,y^{\prime}\in\Gamma}|y^{\prime}\rangle\langle y|\left[\sum_{y^{\prime\prime}\notin\Gamma}\left(\begin{array}[]{cc}U_{y^{\prime\prime}-y}D_{y^{\prime}-y^{\prime\prime}}&0\\ 0&D_{y^{\prime\prime}-y}U_{y^{\prime}-y^{\prime\prime}}\end{array}\right)\right]$ (24)
@@ II.2.2 Scale-dependent topological transitions from EB band engineering
where $\delta\sigma_{x}\otimes\left(T_{-1}+T_{+1}\right)$ represents criss-crossing NNN couplings with strength $\delta$, and $T_{\pm 1}$ denotes the translational operator across one site in the $y$ direction. Remarkably, the size of the $\bar{P}$ unit cell $\left(2y_{R}\times 2y_{R}\right)$ now dramatically influences the effective inter-unit cell coupling strengths and, consequently, the topological phase boundary. As shown in Fig. 3(b-d), changing the unit cell size $2y_{R}=2(L_{y}-1)$ can induce both trivial-to-nontrivial (Fig. 3(b)) and nontrivial-to-trivial (Fig. 3(c)) topological transitions. Below, we shall describe the empirical observation of these transitions, leaving their rigorous derivations and generalizations to the next subsection.
The first scenario with topological trivial-to-nontrivial transition is demonstrated in Fig. 3(c) for illustrative parameters $t_{1}=0.2,t_{2}=0.1,\delta=0.001$. Focusing on the rightmost EB bands (circled), as $L_{y}$ increases from 10 [Fig. 3(b1)] to 42 [Fig. 3(b3)], the bulk EB bands (purple) close up, only to reopen at larger $L_{y}$ with midgap topological EB states (green) [Fig. 3(b4)]. While topological modes (green) also appear at some $L_{y}$ [Fig. 3(b2,b4)] in the second right band, which is not an EB band, their emergence is more erratic with a smaller gap.
Conversely, Fig. 3(d) illustrates the second scenario with a topological nontrivial-to-trivial transition for a different set of parameters $t_{1}=0.05,t_{2}=0.1,\delta=0.001$. Initially topological at small $L_{y}=10$ [Fig. 3(c1)], increasing $L_{y}$ closes up the topological gap, with the critical transition point occurring at approximately $L_{y}=26$ [Fig. 3(c3)], characterized by band touching. Beyond this point, the EB band transitions into a trivial topological phase, as evidenced in Fig. 3(c4).
@@ II.3.1 Construction of generic EB band models
$H^{(\text{EBB})}_{\text{Gen}}$ $=$ $\mathbb{I}\otimes\bar{P}+\sum_{x,Y,\Delta x,\Delta Y}t_{\Delta x,\Delta Y}|x+\Delta x,Y+\Delta Y\rangle\langle x,Y|$
$=$ $\mathbb{I}\otimes\bar{P}+\sum_{k_{x},Y}\sum_{\Delta x,\Delta Y}t_{\Delta x,\Delta Y}e^{ik_{x}\Delta x}T_{\Delta Y}|k_{x},Y\rangle\langle k_{x},Y|,$
@@
To summarize for the generic 1D case, starting from a translation-invariant 1D lattice with hopping amplitudes $t_{\Delta x}$, we can promote it into a 2D system with EB bands by replacing each unit cell by $\bar{P}$. The resultant generalized 2D EB band model is obtained by refining $t_{\Delta x}\rightarrow t_{\Delta x,\Delta Y}$ to include intra-cell hoppings $\Delta Y$, as prescribed by Eq. (LABEL:eq:genhopping).
@@ II.3.3 Scaling dependence in our NNN EB band model
$t_{\Delta x=0,\Delta Y=0}$ $=$ $e^{i\theta}\left(\begin{array}[]{cc}0&t_{1}\\ t_{1}&0\end{array}\right),$ (38)
$t_{\Delta x=-1,\Delta Y=0}$ $=$ $e^{i\theta}\left(\begin{array}[]{cc}0&0\\ t_{2}&0\end{array}\right),$ (41)
$t_{\Delta x=0,\Delta Y=\pm 1}$ $=$ $e^{i\theta}\left(\begin{array}[]{cc}0&\delta\\ \delta&0\end{array}\right).$ (44)
@@
$t_{1}^{\prime}(L_{y})$ $=$ $t_{1}+\delta\,\Omega_{\Delta Y}(L_{y})$
$t_{2}^{\prime}(L_{y})$ $=$ $t_{2}+\delta\,\Omega_{\Delta Y}(L_{y}),$ (46)
@@ II.3.4 Scaling dependence from generic EB band hoppings
Having pinpointed the origin of $L_{y}$-scaling dependence to the $\Omega_{\Delta Y}(L_{y})=\sum_{Y}\tilde{\phi}_{Y+\Delta Y}\phi_{Y}$ term in $H^{\text{Eff}}_{\text{Gen}}$ [Eq. (30) and 31], we now derive its exact $L_{y}$-scaling behavior. To keep the discussion general, we shall not assume any particular form of $P(k)$ or lattice regularization, other than the requirement of geometric defectiveness i.e. that $U(k)=D(k)^{-1}$ diverges at $k\rightarrow 0$ [Eq. (2)].
@@
$\left.\phi_{Y}\right|_{Y=2y-1}$ $=u_{y},\qquad\left.\phi_{Y}\right|_{Y=2y}=v_{y}$
@@ III Discussion
Experimentally, EB band engineering and their scaling-induced transitions can be feasible demonstrated in any non-Hermitian physical setup that can support multiple intra-cell couplings across several or more atoms or orbitals. This include various metamaterial platforms, especially photonic crystals and electrical circuit arrays [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 44, 68, 69], as well as superconducting circuit-based quantum simulators [70, 71, 72, 73, 74, 75]. In particular, the versatility of circuit setups allows for such couplings to be implemented at will. As suggested by our topological phase diagrams, experimental smoking-guns for EB mechanisms can be formulated under diverse ranges of parameters, thanks to the sensitivity and versatility of EB band engineering.
@@ IV Methods
Figure 6: Rotationn of the band spectra of the EB band model of Eq. (53) via $\theta$, under double OBCs. The lattice is composed of $L_{y}=10$ unit cells of $\bar{P}$ in the $y$ direction. The parameters are set as $y_{R}=9$, $a_{0}=10$ and $B=2$. The phase factor $\theta$ is varied from (a) 0 to (b) $\pi/6$, (c) $\pi/4$ and (d) $\pi/3$. The rotation of the spectrum by $\theta$ degrees in the complex plane of ReEE-ImEE is evident. The color ($\chi_{y}$ in Eq. (54)) is used to distinguish EB and non-EB bands.
@@ IV.1 Promoting EB states to EB bands – minimal example
The spectrum of $H_{\text{NN}}$ in Fig. 5(c) consists of highly degenerate middle bands (blue) and isolated non-degenerate bands (red). To verify that they are inherited from non-EB and EB states, we compute their EB projection fraction $\chi_{y}$, which measures the overlap between the $x$-component of an eigenstate $|\Psi(x,y)\rangle$ where $H_{\text{NN}}|\Psi(x,y)\rangle=E_{\text{NN}}|\Psi(x,y)\rangle$ and the EB states $|\phi\rangle$ satisfying $\mathbb{I}_{2}\otimes\bar{P}|\phi\rangle=\bar{p}|\phi\rangle$. Specifically, $\chi_{y}$ is defined as
$\chi_{y}=\left|\langle\phi|\left(\sum_{x}|\Psi_{i}(x,y)\rangle^{2}\right)^{1/2}\right|.$ (54)
@@
Figure 7: Scaling behavior of right and left EB basis states with respect to system size $L_{y}$ and site index $Y$. (a) The calculated amplitudes of right EB states show odd sites (solid curves, a1) and even sites (dashed curves, a2) closely matching $u_{y}$ and $v_{y}$ respectively, across multiple $B$ and system sizes $L_{y}$. (b) The left EB state amplitudes, obtained as eigenvectors of $\bar{P}^{\dagger}$, demonstrate odd sites (solid curves, b1) and even sites (dashed curves, b2) following $\mathcal{N}v_{y}$ and $\mathcal{N}u_{y}$ scaling, respectively, consistently verified for various $B$ and $L_{y}$ values. It is noted that $\mathcal{N}$ [Eq. (62)] is dependent on the system size $L_{y}$ and the specific $B$ value, and is serves to preserve the biorthogonal normalization $\langle\tilde{\phi}|\phi\rangle=1$. By analyzing the scaling behaviors of these right and left EB states, we can obtain the scalings of the effective couplings $t_{1}^{\prime}$ and $t_{2}^{\prime}$ [Eq. (52)] and predict the topological transition points.
@@ IV.2.1 Scaling of EB basis states and $\Omega_{\Delta Y}(L_{y})$ from correlation function behavior
$\left.\phi_{Y}\right|_{Y=2y-1}$ $=u_{y},\qquad\left.\phi_{Y}\right|_{Y=2y}=v_{y}$
@@
For the left EB states, which are eigenvectors of $H^{\dagger}$, their profiles can be directly related to those of the right EB states, as specified by Eq. (57). Essentially, the left ($|\tilde{\phi}\rangle$) and right ($|\phi\rangle$) EB states are spatial inversions of each other, with odd and even site amplitudes interchanged.
The normalization factor $\mathcal{N}$ above is fixed by the biorthogonality condition $\braket{\tilde{\phi}|\phi}=1$, i.e., $\braket{\tilde{\phi}|\phi}=\sum_{Y}\tilde{\phi}_{Y}\phi_{Y}=\sum_{y}\left(\tilde{\phi}_{2y-1}\phi_{2y-1}+\tilde{\phi}_{2y}\phi_{2y}\right)=\sum_{y}2\mathcal{N}u_{y}v_{y}=1$, which gives
@@
$\Omega_{\Delta Y}\left(L_{y}\right)$ $=$ $\sum_{Y}\tilde{\phi}_{Y+\Delta Y}\phi_{Y}$
$=$ $\begin{cases}\mathcal{N}\sum_{y}\left(v_{y+\frac{\Delta Y}{2}}u_{y}+u_{y+\frac{\Delta Y}{2}}v_{y}\right)&\text{even }\Delta Y\quad\\ \mathcal{N}\sum_{y}\left(u_{y+\frac{\Delta Y-1}{2}}u_{y}+v_{y+\frac{\Delta Y+1}{2}}v_{y}\right)&\text{odd }\Delta Y\quad\end{cases}$

### v2

@@ I Introduction
In this work, however, we show that anomalous non-Hermitian scaling can occur without any skin effect at all. Departing from existing paradigms [1, 2, 4, 2, 19, 3, 20, 21, 7, 8, 22], we propose a completely different mechanism for scale-induced topological transitions. Our approach relies on unique renormalization properties near an exceptional point (EP) due to eigenspace defectiveness. While EPs are already known for their extremely sensitive band dispersions and enhanced response in sensing-oriented settings [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34] and unconventional entanglement properties [35, 36, 37, 38, 39, 40, 41, 42], what has not been previously harnessed is the non-locality that emerges from projecting onto their defective eigenspaces.
To employ this non-locality for realizing novel scaling transitions, we specifically dimensionally extend these eigenspaces to form so-called exceptional-bound (EB) bands. These are entire energy bands protected by the presence of momentum-space singularities. As we shall show, EB bands possess unique algebraically decaying profiles that scale strongly with system size, harboring the propensity for scale-sensitive band engineering. Starting from a parent exceptional point, we demonstrate how we can build effective models with precisely scale-tunable couplings. Unlike long-range coupling design approaches that tune hopping profiles link-by-link at fixed size [43, 44], our long-range couplings are uniquely generated once the EP projector kernel and truncation rule are specified, with $L_{y}$ serving as the primary control knob. This enables the same system to be topological at some system sizes but non-topological at other sizes, i.e., exhibiting scaling-controlled topological phase transitions. We develop a formalism for engineering such unprecedented scale-controlled transitions, fundamentally expanding the potential of non-Hermitian mechanisms in non-locally controlling entire band structures, beyond the reach of existing sensing application proposals.
@@ II.1 Unconventional scaling behavior from exceptional point defectiveness
[…]
@@
[…]
@@
$=\frac{1}{2}\left(\begin{array}[]{cc}R_{\Gamma}&-R_{\Gamma}UR_{\Gamma}\\ -R_{\Gamma}DR_{\Gamma}&R_{\Gamma}\end{array}\right).$
@@
$=\frac{1}{2}\left(\begin{array}[]{cc}R_{\Gamma}UR_{\Gamma}DR_{\Gamma}-R_{\Gamma}&0\\ 0&R_{\Gamma}DR_{\Gamma}UR_{\Gamma}-R_{\Gamma}\end{array}\right)$
$=-\frac{1}{2}R_{\Gamma}\left(\begin{array}[]{cc}UR^{c}_{\Gamma}U^{-1}&0\\ 0&U^{-1}R^{c}_{\Gamma}U\end{array}\right)R_{\Gamma}$
$=-\frac{1}{2}\sum_{y,y^{\prime}\in\Gamma}|y^{\prime}\rangle\langle y|\left[\sum_{y^{\prime\prime}\notin\Gamma}\left(\begin{array}[]{cc}U_{y^{\prime\prime}-y}D_{y^{\prime}-y^{\prime\prime}}&0\\ 0&D_{y^{\prime\prime}-y}U_{y^{\prime}-y^{\prime\prime}}\end{array}\right)\right]$
@@ II.2.2 Scale-dependent topological transitions from EB band engineering
where $\delta\sigma_{x}\otimes\left(T_{-1}+T_{+1}\right)$ represents criss-crossing NNN couplings with strength $\delta$, and $T_{\pm 1}$ denotes the translational operator across one site in the $y$ direction. Remarkably, the size of the $\bar{P}$ unit cell $\left(2y_{R}\times 2y_{R}\right)$ now dramatically influences the effective inter-unit cell coupling strengths and, consequently, the topological phase boundary. As shown in Fig. 3(b-d), changing the unit cell size $2y_{R}=2(L_{y}-1)$ can induce both trivial-to-nontrivial (Fig. 3(c)) and nontrivial-to-trivial (Fig. 3(d)) topological transitions. Below, we shall describe the empirical observation of these transitions, leaving their rigorous derivations and generalizations to the next subsection.
The first scenario with topological trivial-to-nontrivial transition is demonstrated in Fig. 3(c) for illustrative parameters $t_{1}=0.2,t_{2}=0.1,\delta=0.001$. Focusing on the rightmost EB bands (circled), as $L_{y}$ increases from 10 [Fig. 3(c1)] to 42 [Fig. 3(c3)], the bulk EB bands (purple) close up, only to reopen at larger $L_{y}$ with midgap topological EB states (green) [Fig. 3(c4)]. While topological modes (green) also appear at some $L_{y}$ [Fig. 3(c2,c4)] in the second right band, which is not an EB band, their emergence is more erratic with a smaller gap.
Conversely, Fig. 3(d) illustrates the second scenario with a topological nontrivial-to-trivial transition for a different set of parameters $t_{1}=0.05,t_{2}=0.1,\delta=0.001$. Initially topological at small $L_{y}=10$ [Fig. 3(d1)], increasing $L_{y}$ closes up the topological gap, with the critical transition point occurring at approximately $L_{y}=26$ [Fig. 3(d3)], characterized by band touching. Beyond this point, the EB band transitions into a trivial topological phase, as evidenced in Fig. 3(d4).
@@ II.3.1 Construction of generic EB band models
$H^{(\text{EBB})}_{\text{Gen}}$ $=\mathbb{I}\otimes\bar{P}+\sum_{x,Y,\Delta x,\Delta Y}t_{\Delta x,\Delta Y}|x+\Delta x,Y+\Delta Y\rangle\langle x,Y|$ (12)
$=\mathbb{I}\otimes\bar{P}+\sum_{k_{x},Y}\sum_{\Delta x,\Delta Y}t_{\Delta x,\Delta Y}e^{ik_{x}\Delta x}T_{\Delta Y}|k_{x},Y\rangle\langle k_{x},Y|.$
@@
To summarize for the generic 1D case, starting from a translation-invariant 1D lattice with hopping amplitudes $t_{\Delta x}$, we can promote it into a 2D system with EB bands by replacing each unit cell by $\bar{P}$. The resultant generalized 2D EB band model is obtained by refining $t_{\Delta x}\rightarrow t_{\Delta x,\Delta Y}$ to include intra-cell hoppings $\Delta Y$, as prescribed by Eq. (12).
@@ II.3.3 Scaling dependence in our NNN EB band model
$t_{\Delta x=0,\Delta Y=0}$ $=e^{i\theta}\left(\begin{array}[]{cc}0&t_{1}\\ t_{1}&0\end{array}\right),$
$t_{\Delta x=-1,\Delta Y=0}$ $=e^{i\theta}\left(\begin{array}[]{cc}0&0\\ t_{2}&0\end{array}\right),$
$t_{\Delta x=0,\Delta Y=\pm 1}$ $=e^{i\theta}\left(\begin{array}[]{cc}0&\delta\\ \delta&0\end{array}\right).$
@@
$t_{1}^{\prime}(L_{y})$ $=t_{1}+\delta\,\Omega_{\Delta Y}(L_{y}),$ (18)
$t_{2}^{\prime}(L_{y})$ $=t_{2}+\delta\,\Omega_{\Delta Y}(L_{y}).$
@@ II.3.4 Scaling dependence from generic EB band hoppings
Having pinpointed the origin of $L_{y}$-scaling dependence to the $\Omega_{\Delta Y}(L_{y})=\sum_{Y}\tilde{\phi}_{Y+\Delta Y}\phi_{Y}$ term in $H^{\text{Eff}}_{\text{Gen}}$ [Eq. (13) and 14], we now derive its exact $L_{y}$-scaling behavior. To keep the discussion general, we shall not assume any particular form of $P(k)$ or lattice regularization, other than the requirement of geometric defectiveness i.e. that $U(k)=D(k)^{-1}$ diverges at $k\rightarrow 0$ [Eq. (2)].
@@
$\left.\phi_{Y}\right|_{Y=2y-1}$ $=u_{y},\qquad\left.\phi_{Y}\right|_{Y=2y}=v_{y}$ (20)
@@ III Discussion
Experimentally, EB band engineering and their scaling-induced transitions can be feasible demonstrated in any non-Hermitian physical setup that can support multiple intra-cell couplings across several or more atoms or orbitals. This include various metamaterial platforms, especially photonic crystals and electrical circuit arrays [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 48, 72, 73, 74, 75], as well as superconducting circuit-based quantum simulators [76, 77, 78, 79, 80]. In particular, the versatility of circuit setups allows for such couplings to be implemented at will. As suggested by our topological phase diagrams, experimental smoking-guns for EB mechanisms can be formulated under diverse ranges of parameters, thanks to the sensitivity and versatility of EB band engineering.
@@ IV Methods
Figure 6: Rotation of the band spectra of the EB band model of Eq. (26) via $\theta$, under double OBCs. The lattice is composed of $L_{y}=10$ unit cells of $\bar{P}$ in the $y$ direction. The parameters are set as $y_{R}=9$, $a_{0}=10$ and $B=2$. The phase factor $\theta$ is varied from (a) 0 to (b) $\pi/6$, (c) $\pi/4$ and (d) $\pi/3$. The rotation of the spectrum by $\theta$ degrees in the complex plane of ReEE-ImEE is evident. The color ($\chi_{y}$ in Eq. (27)) is used to distinguish EB and non-EB bands.
@@ IV.1 Promoting EB states to EB bands – minimal example
[…]
@@
[…]
@@ IV.2.1 Scaling of EB basis states and $\Omega_{\Delta Y}(L_{y})$ from correlation function behavior
$\left.\phi_{Y}\right|_{Y=2y-1}$ $=u_{y},\qquad\left.\phi_{Y}\right|_{Y=2y}=v_{y}$ (30)
@@
[…]
The normalization factor $\mathcal{N}$ above is fixed by the biorthogonality condition $\langle\tilde{\phi}|\phi\rangle=1$, i.e., $\langle\tilde{\phi}|\phi\rangle=\sum_{Y}\tilde{\phi}_{Y}\phi_{Y}=\sum_{y}\left(\tilde{\phi}_{2y-1}\phi_{2y-1}+\tilde{\phi}_{2y}\phi_{2y}\right)=\sum_{y}2\mathcal{N}u_{y}v_{y}=1$, which gives
@@
$\Omega_{\Delta Y}\left(L_{y}\right)$ $=\sum_{Y}\tilde{\phi}_{Y+\Delta Y}\phi_{Y}$ (37)
[…]